# Separable models for dynamic signed networks

### v1

@@ 8.1 Software
The new B2Lstergm package for R (R Development Core Team, 2025), developed alongside this paper, represents a significant extension of existing tools for Bayesian inference on temporal network models. Specifically designed for Bayesian separable 2-layer TERGMs on signed networks, B2Lstergm is currently under active development and aims to offer a more flexible and computationally efficient framework than what is currently available. It builds on and extends the modelling and computational capabilities of the Bergm package (Caimo et al., 2022, 2025), which implements Bayesian inference for static ERGMs. Additionally, the multi.ergm package (Krivitsky, 2025), which supports multi-layer ERGMs (Krivitsky, Koehly and Marcum, 2020), is employed here to simulate network statistics corresponding to the interaction and conditional sign processes of the two separable likelihood components described in Equation 6. This simulation relies on the efficient MCMC algorithms implemented in the ergm package (Hunter et al., 2008; Krivitsky et al., 2023; Handcock et al., 2025). Notably, although the stergm() function within the tergm package (Krivitsky and Handcock, 2025) typically relies on stochastic maximum likelihood estimation, it can alternatively use contrastive divergence when computational efficiency is a priority or convergence issues arise.
@@ 9 Application to US Congressional data
We employ our model to examine the evolution of political cooperation and opposition in the United States Congress using the signed network data inferred by Neal (2020, 2014) from bill co-sponsorship data via stochastic degree sequence model.
Our focus is on the 99th-101st Congresses (1985-1991), characterised by a hung government with a split majority. This period saw a republican president, Ronal Reagan, governing alongside a Democratic majority in the Senate. An examination of the Senate’s behaviour during this time can provide valuable insights into the dynamics of the Senate and its role in shaping policy during a period of significant upheaval (Rohde, 1991). This period encompasses several pivotal events, including the collapse of the Soviet Union (1989-1991), the Gulf War (1990-1991), and the Savings and Loan Crisis (1989-1991).
@@
Figure 5: US Senators signed network. Blue and red nodes represent Democratic and Republican Senators, respectively. Graphs: (a) signed relationships during the 99th Congress (initial state); (b) signed relationships during the 100th Congress; (c) signed relationships during the 101th Congress.
@@ 9.1 Model specification
We adopt a time-homogeneous separable 2-layer TERGM where the formation and persistence configurations are the same and the list of statistics used is specified below. Obviously, in the case of excessive non-stationarity across multiple network observations, the use of a homogeneous model could prove problematic.
@@
The interaction network exhibits such high density that it lacks discernible structural heterogeneity. Consequently, we restrict our analysis to the conditional signed process and the posterior distribution of its associated parameters, $\pi(\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}}\mid\mathbf{z}_{0:T},\mathbf{x}_{0:T})$. We define a separable 2-layer TERGM using the following network statistics:
@@
    number of positive edges, ${}^{+},$ measuring the density of positive interactions;
@@
We assign independent, weakly informative normal priors to the model parameters, informed in part by the structural characteristics of the initial network $\mathbf{y}_{0}$. Specifically, we set the prior means for both $\zeta^{\mathcal{F}}_{\text{edges}^{+}}$ and $\zeta^{\mathcal{P}}_{\text{edges}^{+}}$ to –1, and the prior mean for $\zeta^{\mathcal{F}}_{\text{gwesf}^{+}}(\alpha)$ is set accordingly. All prior variances are fixed at 25.
@@ 9.2 Posterior inference
Posterior inference is conducted using the adaptive Metropolis–Hastings approximate exchange algorithm (MH-AEA), as outlined in Algorithm (1). Each iteration employs $5000$ auxiliary draws for the network simulation, and the main Markov chain comprises $35000$ iterations, with the initial $10000$ discarded as burn-in. The MCMC trace plots (Figure 6) of simulated posterior parameters indicate that the MCMC chain is convergent to the stationary posterior distribution. Table 2 presents a summary of the marginal posterior parameter density estimates.
@@
Posterior predictive checks are performed to evaluate how well the model captures the observed data. Simulated network statistics were generated from the posterior predictive distribution and compared to the observed statistics using boxplots (Figure 7). In these plots, the medians of the posterior predictive statistics are close to $0,$ indicating that the observed data are consistent with what the model predicts. This suggests that the model provides a good fit to the data in terms of model statistics.
@@
Table 2: Summary of the posterior parameter density estimates.
@@
Parameters Mean 0.025 0.5 0.975 Mean 0.025 0.5 0.975
$\zeta_{\texttt{edges}^{+}}$3.95 -5.81 -3.98 -2.00 -4.77 -5.90 -4.76 -3.68
@@
$\zeta_{\texttt{gwdegree}^{+}(\alpha=0.2)}$0.29 -3.11 -0.31 2.49 0.81 -0.28 0.79 1.87
@@
$\zeta_{\texttt{gwese}^{+}(\alpha=0.6)}$1.41 -2.11 -1.40 -0.70 -0.58 -0.96 -0.58 -0.20
$\zeta_{\texttt{gwese}^{-}(\alpha=0.6)}$1.05 -1.41 -1.03 -0.75 -0.62 -0.81 -0.61 -0.48
@@
  • The strongly negative 95% posterior credible intervals for $\zeta_{\texttt{edges}^{+}}$ in both the formation and persistence processes suggest that the presence of an existing positive edge reduces the likelihood of forming or maintaining additional positive interactions. This may reflect either competition for alliances or strategic repositioning in response to institutional instability.
  • The 95% posterior credible interval for $\zeta_{\texttt{homophily}^{+}(\texttt{rep})}$ is slightly positive for formation but not significantly different from 0 for persistence, indicating that while Republican senators exhibit a modest tendency to form positive edges within the party, these interactions are not particularly stable over time. This instability may point to underlying intra-party conflict.
  • The estimate for $\zeta_{\texttt{gwdegree}^{+}(\alpha=0.2)}$ in both processes is not significantly different from zero, suggesting a relatively even distribution of interactions, implying no strong centralisation in the network structure.
  • The consistently positive 95% posterior credible intervals for $\zeta_{\texttt{gwesf}^{+}(\alpha=0.6)}$ provide robust support for the role of balanced triadic closure in both initiating and sustaining positive interactions. This is consistent with the theoretical expectation of structural balance and indicates a preference for cohesive, reinforcing social triangles.
  • The negative posterior estimates for $\zeta_{\texttt{gwese}^{+}(\alpha=0.6)}$ and $\zeta_{\texttt{gwese}^{-}(\alpha=0.6)}$ (also negative across both processes) suggest that unbalanced triadic structures – whether involving positive or negative edges – are systematically avoided. This aligns with strong structural balance theory.
Collectively, these findings suggest a polarised but strategically coherent interaction network, in which alliances remain fluid and structural balance emerges as a more stable organising mechanism than party affiliation.

### v2

@@ 8.1 Software
The B2Lstergm package for R (R Development Core Team, 2025), developed alongside this paper, represents a significant extension of existing tools for Bayesian inference on temporal network models. It builds on and extends the modelling and computational capabilities of the Bergm package (Caimo et al., 2022, 2025), which implements Bayesian inference for static ERGMs. Additionally, the multi.ergm package (Krivitsky, 2025), which supports multi-layer ERGMs (Krivitsky et al., 2020), is employed here to simulate network statistics corresponding to the interaction and conditional sign processes of the two separable likelihood components described in Equation 5. This simulation relies on the efficient MCMC algorithms implemented in the ergm package (Hunter et al., 2008; Krivitsky et al., 2023; Handcock et al., 2025). Notably, although the tergm() function within the tergm package (Krivitsky and Handcock, 2025) typically relies on stochastic maximum likelihood estimation (MLE), it can alternatively use contrastive divergence when computational efficiency is a priority or convergence issues arise. The current implementation of the B2Lstergm package can handle networks with a few hundred interacting edges. For larger networks, however, more efficient approximate methods, such as the pseudo-posterior adjustment proposed in Bouranis et al. (2017), would be necessary.
9 Comparative simulation study
Among existing approaches, only the TERGM by Fritz et al. (2025) and our approach directly extend the ERGM framework to signed networks. In this simulation study, we compare three modelling approaches with the aim of examining how their interpretative properties differ.
We consider the TERGM proposed by Fritz et al. (2025), described in Section 5, as well as our 2-layer STERGM. In addition, we introduce an intermediate modelling approach: a separable temporal extension of the TERGM in Fritz et al. (2025). In this STERGM formulation, the transition probability is defined as
$p\left(\mathbf{y}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\vartheta}^{\mathcal{F}},\boldsymbol{\vartheta}^{\mathcal{P}}\right)=p\left(\mathbf{y}^{\mathcal{F}}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\vartheta}^{\mathcal{F}}\right)\times p\left(\mathbf{y}^{\mathcal{P}}_{t}\mid\mathbf{y}_{t-1},\boldsymbol{\vartheta}^{\mathcal{P}}\right).$
It is important to note that when $\mathbf{x}$ is complete, i.e., all nodes interact with all others, the STERGM and the 2-layer STERGM coincide, as $\mathbf{y}$ is binary, encoding the sign of each edge, and therefore equivalent to $\mathbf{z}.$
To better illustrate the interpretative differences between the three approaches, we focus on a simple specification including density and positive transitivity terms. Specifically, we generate a random graph on $40$ nodes for the initial interaction network $\mathbf{x}_{1}$, where each interaction occurs with probability $0.2$, and each realised edge is independently assigned a positive or negative sign with probability $0.5$. The interaction network $\mathbf{x}_{2}$ is then generated using a TERGM, with both the baseline edge density and the change-of-state density parameters set to 2-2. Conditional on the transition from $\mathbf{x}_{1}$ to $\mathbf{x}_{2}$ we simulate a 2-layer STERGM to generate $\mathbf{z}_{2}$ specified as follows:
${\zeta_{\texttt{edges}^{+}}^{\mathcal{F}}}=0,\qquad{\zeta_{\texttt{gwesf}^{+}}^{\mathcal{F}}}(\alpha)=0.2,\qquad{\zeta_{\texttt{edges}^{+}}^{\mathcal{P}}}=0,\qquad{\zeta_{\texttt{gwesf}^{+}}^{\mathcal{P}}}(\alpha)=0.$
Here, $\texttt{edges}^{+}$ denotes the number of positive edges in the formation and dissolution processes, and $\texttt{gwesf}^{+}$ is the geometrically weighted positive edgewise shared friends (see Figure 4) representing the statistics related to positive triadic closure. We set the decay parameter $\alpha$ for the $\texttt{gwesf}^{+}$ terms to $0.6$ In practice, the only non-zero effect is therefore the formation parameter $\zeta_{\texttt{gwesf}^{+}}^{\mathcal{F}}$}, which encourages the creation of positive triangles from $t_{1}$ to $t_{2}$.
The TERGM approach includes terms for the stability of positive edges, changes in positive edges, and changes in transitive triads. The STERGM version of Fritz et al. (2025) uses the same type of statistics employed in the simulation, as does the 2-layer STERGM.
In Figure 5, we can observe the MLE estimates and their corresponding 95% confidence intervals for both the TERGM and STERGM approaches, computed using the ergm.sign package for R (Schalberger et al., 2025). We also show the posterior means and 95% credible intervals for the two-layer STERGM, obtained with the B2Lstergm package using flat priors on all parameters, in order to make the comparison easier.
Figure 5: MLE estimates (black dots) with 95% confidence intervals are shown for the TERGM and STERGM based on Fritz et al. (2025), while the 2-layer STERGM displays posterior means (black dots) with 95% credible intervals. In the 2-layer STERGM plot, the white dots with black outline indicate the true parameter values used to generate the data.
We can see that the positive and negative density parameters in the TERGM correctly capture the drop in overall density caused by the reduced level of interaction from $t_{1}$ to $t_{2}$. The parameter $\vartheta_{\texttt{gwesf}^{+}}$, associated with positive triangles, is not significantly different from zero. In this model, without additional controls for other forms of transitive structure, it is not possible to determine whether this value reflects a real counteracting effect, or whether the change instead arises from substantially greater formation or proportionally less dissolution of transitive structures, relative to the others. In practice, this model determines dyad change rates according to the observed probability of each dyad switching state. Therefore, when the initial network has a density below $0.5$, there are more opportunities for formation, whereas a density above $0.5$ results in more opportunities for dissolution.
In the STERGM, the parameter $\vartheta_{\texttt{edges}^{+}}^{\mathcal{F}}$} is negative, as it reflects the fact that the proportion of positive edges formed between $t_{1}$ and $t_{2}$ is negative due to the non significant increase in the proportion of positive edges in the formation process (relative to the proportion of negative edges and empty dyads). The parameter $\vartheta_{\texttt{gwesf}^{+}}^{\mathcal{F}}$} is not significant. If we add a control for transitive interaction in the formation process, $\vartheta_{\texttt{gwesf}^{+}}^{\mathcal{F}}$} becomes significantly positive ($5.46$ with a standard error of about $0.17$), balancing a significantly negative parameter (5.65-5.65 with a standard error of about $0.16$). We can therefore conclude that this categorical model captures the increase in positive transitive structures, but we cannot determine precisely whether this is due to the formation of genuinely new positive interactions or to the transformation of existing non-positive configurations into positive ones. In other words, we cannot clearly separate the effect of sign changes from the formation interaction process itself.
The 2L-STERGM model, through the parameter $\zeta_{\texttt{gwesf}^{+}}^{\mathcal{F}}$}, captures the increase in the proportion of positive transitive relations at the expense of other types of triadic signed configurations, allowing us to deduce that there has been a rise in positive transitive structures, net of overall structural changes in the interaction network. The greater uncertainty in the persistence process, arises from the reduced sample space determined by the low number of persistent interactions from $t_{1}$ to $t_{2}$.
@@ 10 Application to US Congressional data
We employ our model to examine the evolution of political cooperation and opposition in the United States Congress using the signed network data inferred by Neal (2014, 2020) from bill co-sponsorship data via stochastic degree sequence model (SDSM).
Our focus is on the 99th-101st Congresses (1985-1991), characterised by a hung government with a split majority. This period saw a republican president, Ronal Reagan, governing alongside a Democratic majority in the Senate. An examination of the Senate behaviour during this time can provide valuable insights into the dynamics of the Senate and its role in shaping policy during a period of significant upheaval (Rohde, 1991). This period encompasses several pivotal events, including the collapse of the Soviet Union (1989-1991), the Gulf War (1990-1991), and the Savings and Loan Crisis (1989-1991).
When Congressional co-sponsorship data are modelled using the SDSM framework, each legislator has expected positive and negative degree tendencies. For each dyad, the probability of a positive, negative, or no edge is computed based on the product of their respective tendencies. Edges are then assigned probabilistically: dyads with higher positive tendencies are more likely to become positive edges, those with higher negative tendencies are more likely to become negative edges, and others remain unconnected. This produces a signed network that reflects legislators’ expected co-sponsorship and avoidance patterns.
@@
Figure 6: US Senators signed network. Blue and red nodes represent Democratic and Republican Senators, respectively. Black and orange edges represent positive and negative relationships, respectively. Graphs: (a) signed relationships during the 99th Congress (initial state); (b) signed relationships during the 100th Congress; (c) signed relationships during the 101st Congress.
@@ 10.1 Model specification
We adopt a time-homogeneous 2-layer STERGM where the formation and persistence configurations are the same and the list of statistics used is specified below. Obviously, in the case of excessive non-stationarity across multiple network observations, the use of a homogeneous model could prove problematic.
@@
The interaction network exhibits such high density that it lacks discernible structural heterogeneity. Consequently, we restrict our analysis to the conditional signed process and the posterior distribution of its associated parameters $\pi(\boldsymbol{\zeta}^{\mathcal{F}},\boldsymbol{\zeta}^{\mathcal{P}}\mid\mathbf{z}_{0:T},\mathbf{x}_{0:T})$. We define a 2-layer STERGM using the following network statistics:
@@
[…]
@@
We assign independent, weakly informative Normal priors to all model parameters, incorporating mild guidance from the structural characteristics of the initial network. Specifically, the prior means for both $\zeta^{\mathcal{F}}_{\text{edges}^{+}}$ and $\zeta^{\mathcal{P}}_{\text{edges}^{+}}$ are set to 1–1, reflecting a prior belief that the baseline probability of a positive edge (not involved in triadic or degree-based configurations) is low, but without imposing a strong constraint on its magnitude. Similarly, the prior mean for $\zeta^{\mathcal{F}}_{\text{gwesf}^{+}}(\alpha)$ is set to reflect moderate triadic closure, consistent with expectations of structural balance in cooperative interactions. All priors are assigned a large variance of 25, ensuring they remain weakly informative, vague enough to let the data dominate inference, while still preventing numerical instabilities and implausible parameter values during MCMC estimation.
@@ 10.2 Posterior inference
Posterior inference is conducted using the adaptive AEA, as outlined in Algorithm (1). Each iteration employs $5000$ auxiliary draws for the network simulation, and the main Markov chain comprises $35000$ iterations, with the initial $10000$ discarded as burn-in. MCMC trace plots (Figure 7) of the simulated posterior parameters indicate that the MCMC chain converges to the stationary posterior distribution. Table 2 presents a summary of the posterior density parameter estimates.
@@
The quality of the posterior estimates was evaluated using autocorrelation plots (Figure 8) and effective sample sizes. All parameters exhibited effective sample sizes greater than 200, suggesting adequate chain convergence and mixing.
Figure 8: MCMC autocorrelation plots for each parameter.
Posterior predictive checks are performed to assess how well the model captures the observed data. Simulated network statistics were generated from the posterior predictive distribution and compared to the observed statistics using boxplots (Figure 9). In these plots, the medians of the posterior predictive statistics are close to $0,$ indicating that the observed data are consistent with what the model predicts. This suggests that the model provides a good fit to the data in terms of model statistics.
@@
Table 2: Summary of the posterior parameter density estimates of the conditional sign process.
@@
Parameters Mean 2.5% Median 97.5% Mean 2.5% Median 97.5%
$\zeta_{\texttt{edges}^{+}}$ -3.95 -5.81 -3.98 -2.00 -4.77 -5.90 -4.76 -3.68
@@
$\zeta_{\texttt{gwdegree}^{+}(\alpha=0.2)}$ -0.29 -3.11 -0.31 2.49 0.81 -0.28 0.79 1.87
@@
$\zeta_{\texttt{gwese}^{+}(\alpha=0.6)}$ -1.41 -2.11 -1.40 -0.70 -0.58 -0.96 -0.58 -0.20
$\zeta_{\texttt{gwese}^{-}(\alpha=0.6)}$ -1.05 -1.41 -1.03 -0.75 -0.62 -0.81 -0.61 -0.48
@@
  • The strongly negative 95% posterior credible intervals for $\zeta_{\texttt{edges}^{+}}$ in both the formation and persistence processes indicate that the baseline probability of forming or maintaining a positive edge is very low. For example, the posterior mean of $\zeta_{\texttt{edges}^{+}}=-3.95$ in the formation process corresponds to a conditional probability of forming a positive edge given an interaction of $\operatorname{logit}^{-1}(-3.95)\approx 0.019$, and for persistence, $\zeta_{\texttt{edges}^{+}}=-4.77$ gives $\operatorname{logit}^{-1}(-4.77)\approx 0.008.$
  • The 95% posterior credible interval for $\zeta_{\texttt{homophily}^{+}(\texttt{rep})}$ is slightly positive for formation but not significantly different from 0 for persistence. Marginally, this translates to a modest increase in the probability of forming a positive edge among republican senators, e.g., $\operatorname{logit}^{-1}(-3.95+0.53)\approx 0.027$, while the effect on persistence remains negligible.
  • The estimates for $\zeta_{\texttt{gwdegree}^{+}(\alpha=0.2)}$ are not significantly different from zero in either process, implying that high-degree nodes do not systematically form or maintain more positive edges. This suggests a relatively even distribution of interactions across the network.
  • The consistently positive 95% posterior credible intervals for $\zeta_{\texttt{gwesf}^{+}(\alpha=0.6)}$ support the role of triadic closure in both forming and sustaining positive interactions. For instance, adding one unit of balanced triadic closure increases the conditional probability of positive formation from $\approx 0.019$ to $\operatorname{logit}^{-1}(-3.95+2.42)\approx 0.18$ and the probability of persistence from $\approx 0.008$ to $\operatorname{logit}^{-1}(-4.77+1.56)\approx 0.039$.
  • The negative posterior estimates for $\zeta_{\texttt{gwese}^{+}(\alpha=0.6)}$ and $\zeta_{\texttt{gwese}^{-}(\alpha=0.6)}$, where the former corresponds to a balanced triadic configuration and the latter to an unbalanced one, indicate that these structures reduce the likelihood of forming or maintaining positive edges. For example, a dyad involved in an unbalanced negative triad would have its conditional formation probability reduced from $\approx 0.019$ to $\operatorname{logit}^{-1}(-3.95-1.05)\approx 0.0067.$
10.3 Marginal estimation
The network interaction structure $\mathbf{x}_{0:T}$ was analysed by estimating the following STERGM posterior
$\pi\left(\boldsymbol{\xi}^{\mathcal{F}},\boldsymbol{\xi}^{\mathcal{P}}\mid\mathbf{x}_{0:T}\right)\propto\prod_{t=1}^{T}\exp\left\{{\boldsymbol{\xi}^{\mathcal{F}}}^{\top}\mathbf{s}\left(\mathbf{x}_{t}^{\mathcal{F}};\mathbf{y}_{t-1}\right)\right\}\;\exp\left\{{\boldsymbol{\xi}^{\mathcal{P}}}^{\top}\mathbf{s}\left(\mathbf{x}_{t}^{\mathcal{P}};\mathbf{y}_{t-1}\right)\right\}\;\pi\left(\boldsymbol{\xi}^{\mathcal{F}}\right)\;\pi\left(\boldsymbol{\xi}^{\mathcal{P}}\right)$
using a similar version of Algorithm 1. For both the formation and persistence, we used the following three network statistics that are matching the configurations used in the conditional signed process: (1) edges, the baseline propensity for any interaction to form or persist; (2) $\texttt{homophily(rep)},$ the propensity for interactions between republican Senators to form or persist; (3) $\texttt{gwesp}(\alpha=0.2),$ the tendency for interactions to form or persist in closed triads; and (4) $\texttt{gwdegree}(\alpha=0.6),$ the influence of node degree on the formation and persistence of edges. Posterior sampling was performed using a prior specification and initialisation procedure similar to that described in Section 10.2.
Table 3: Summary of the posterior parameter density estimates of the interaction process.
Formation ($\mathcal{F}$) Persistence ($\mathcal{P}$)
Parameters Mean 2.5% Median 97.5% Mean 2.5% Median 97.5%
$\xi_{\texttt{edges}}$ -6.40 -7.45 -6.38 -5.46 -4.78 -5.46 -4.80 -4.12
$\xi_{\texttt{homophily}(\texttt{rep})}$ 0.08 -0.15 0.08 0.31 0.06 -0.30 0.06 0.43
$\xi_{\texttt{gwdegree}(\alpha=0.2)}$ -1.18 -3.65 -1.11 0.89 -0.10 -0.94 -0.08 0.66
$\xi_{\texttt{gwesp}(\alpha=0.6)}$ 2.52 2.04 2.51 3.07 1.55 1.24 1.56 1.88
The results summarised in Table 3 indicate that new interaction formation was relatively infrequent, with a strongly negative baseline tendency ($\xi^{\mathcal{F}}_{\texttt{edges}}$), suggesting that in the absence of other effects, new edges were unlikely to form. There was only weak evidence of Republican homophily in edge formation ($\xi^{\mathcal{F}}_{\texttt{homophily}(\texttt{rep})}$), indicating minimal preference for within-group connections. Actors with higher existing degrees were somewhat less likely to create additional edges ($\xi^{\mathcal{F}}_{\texttt{gwdegree}(\alpha)}$) consistent with degree saturation effects limiting new edge formation. By contrast, there was strong and credible evidence of transitive closure in the formation process ($\xi^{\mathcal{F}}_{\texttt{gwesp}(\alpha)}$), indicating that new edges tend to form within cohesive triads.
In contrast, existing interactions display a strong tendency to persist, although overall edge stability remained selective ($\xi^{\mathcal{P}}_{\texttt{edges}}$). Homophily again played little role in persistence ($\xi^{\mathcal{P}}_{\texttt{homophily}(\texttt{rep})}$). Degree effects were near zero ($\xi^{\mathcal{P}}_{\texttt{gwdegree}(\alpha)}$), implying that high-degree individuals were only marginally less likely to sustain their existing edges. Transitive closure effects were positive ($\xi^{\mathcal{P}}_{\texttt{gwesp}(\alpha)}$), showing that edges embedded in triadic structures were substantially more stable over time.
Overall, these results suggest a relatively stable network characterised by limited new edge creation, strong clustering tendencies within triads, and modest degree-related constraints on both the formation and maintenance of interactions. New connections tend to emerge within existing cohesive subgroups, while established relationships persist preferentially within these clustered structures.
To obtain marginal probabilities for positive edges, we multiply the interaction probabilities by the conditional probabilities according to Equation (4).
The estimated probabilities suggest that the formation of new interactions in the network is rare. The strongly negative formation parameter $\xi_{\texttt{edges}}$ implies that, in the absence of additional structural effects, the probability of any new edge emerging is extremely low, $\text{logit}^{-1}(-6.40)\approx 0.0017$. When this is combined with the small conditional probability that a newly formed interaction is positive ($\approx 0.019$), the resulting marginal probability of observing a new positive edge becomes effectively negligible. However, this pattern changes markedly once triadic closure is considered. Incorporating the transitivity effect $\xi_{\texttt{gwesp}(\alpha=0.6)}$ increases the probability of forming an interaction to about $\operatorname{logit}^{-1}(-6.40+2.52)\approx 0.02$, and the probability that such an interaction is positive to roughly $\operatorname{logit}^{-1}(-3.95+2.42)\approx 0.18$. The resulting marginal probability of forming a positive edge within a triad ($\zeta_{\texttt{gwesf}^{+}}$) rises to around $0.18\times\text{logit}^{-1}(-6.40+2.52)\approx 0.0036$. This clearly indicates that the formation of positive edges is concentrated within cohesive positive triadic structures.
Persistence follows a similar but slightly stronger pattern. Existing interactions are more likely to continue than new ones are to form, with a baseline persistence probability of around $0.0083$. Yet, the probability that a persisting edge remains positive is very close to 0. When triadic effects are present, both persistence and conditional positivity rise to approximately $\text{logit}^{-1}(-4.78+1.55)\approx 0.038$, resulting in a marginal positive persistence of about $0.038\times\text{logit}^{-1}(-4.77+1.56)\approx 0.0015$. These findings imply that while the overall network formation structure remains sparse, positive relationships, though rare, are substantially more likely to both form and endure within tightly clustered triads. Taken together, the results portray a network characterised by isolated sets of stable, mutually reinforcing positive interactions.
While these estimates provide a useful illustration of the patterns in the network, the true uncertainty around these marginal probabilities may be larger, and credible intervals derived from posterior draws provide a more complete picture.